Suppose the monthly income of an individual increases from Rs. 10,000 to Rs. 15,000
which increases his demand for clothes from 20 units to 25 units. Calculate the income
elasticity of demand and interpret the result.
1
Expert's answer
2019-09-23T08:59:44-0400

Income elasticity of demand is the responsiveness of quantity demanded for clothes to change in income of the consumer.

Given,

change in income ("\Delta"I) = 5000 (15000-10000)

and

change in quantity demanded of clothes ("\Delta" Qd) = 5 (25-20)

Initial income ( I) = 10000

Initial quantity demanded of clothes (Qd) = 20


Income elasticity of demand is calculated as:

("\Delta" Qd/Qd )/ ("\Delta"I/I)= (5/20)/(5000/10000)

= (1/4)/ (1/2)

IED = 0.5